The ratio of Fatima’s income to her savings is 4:1. The percentage of money saved by her is ______.

Grade:10

1 Answers

Forum 4
one year ago
Given, ratio of income to savings=4:1
Let income=4xand savings=x.
∴Percentage of money saved by her=Savings/(Income)×100%[We find percentage of saving on the basis of income ]
=(x/4x)×100=(x/4x)×100=1/4×100%=25%

Hope this answer is helpful.